OpenStudy (anonymous):

Give a parameter interval that traces the curve exactly once. x = 3sin(2t) y = 1.5cos(t)

OpenStudy (anonymous):

Yo @jim_thompson5910 buddy can you help me out here a bit? :D

OpenStudy (anonymous):

Start by letting \(t=0\).

OpenStudy (anonymous):

Find the initial position on the interval.

OpenStudy (anonymous):

you get \((x,y)=(0,1.5)\)

OpenStudy (anonymous):

Now we need to find solutions:\[ 0 = 3\sin(2t),\quad 1.5=1.5\cos(t) \]

OpenStudy (anonymous):

Solutions for the first equation are: \[ \frac \pi 2 n \]And solutions for the second one are: \[ 2\pi n \]

OpenStudy (anonymous):

The least common multiple is \(2\pi\) so I'd say that is the period.